OpenStudy (anonymous):

Two numbers add to 436 and the first is 134 bigger than the second. What are the two numbers?

OpenStudy (anonymous):

First set two variables. The second number could be x and the second one could be x+134. If u still don't get it tell me

OpenStudy (scorcher219396):

First number: X+134 ("134 bigger than the second") Second number: X X+(X+134)=436 ("two numbers add to 436")

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

ok collegcountry guy. Lets pretend the second number is x

OpenStudy (anonymous):

because we don;t know what the number is we have to set the second number as x, then we learn that, "first is 134 bigger than the second" so the second number is x+134.

OpenStudy (anonymous):

x+134+x=436

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do u get it know, if u need further help feel free to ask me

OpenStudy (anonymous):

now?

OpenStudy (anonymous):

I'm sorry I don't understand

OpenStudy (anonymous):

could u help me by telling me what u don't get

OpenStudy (anonymous):

the point of using variables

OpenStudy (anonymous):

well we don't know any of the numbers, and variables can just stay in place for 1 of the numbers. So we have x for that number

OpenStudy (anonymous):

Do u get this part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And the question tells us, " the first is 134 bigger than the second. ?"

OpenStudy (anonymous):

So lets say x is the second number.

OpenStudy (anonymous):

Do u get this part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and so the first number is x+134. And the second number would be just x. So our equation would be x+134+x=436

OpenStudy (anonymous):

DO u get this part.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now just solve for x

OpenStudy (anonymous):

TELl me when u got x

OpenStudy (anonymous):

huh? I'm sorry I don't know

OpenStudy (anonymous):

Oh(sry, I am going into 7th grade Algebra 1) x+134+x=436. WEll we would first add the two x's. so we have 2x+134=436. Now we subtract 134 fromm both sides to have x on one side only. So we get 2x=302. Divide 2 on both sides and we get x+151

OpenStudy (anonymous):

Tell me if u get this

OpenStudy (anonymous):

oh i got it now

OpenStudy (anonymous):

Could u tell me what I could have improved on.

OpenStudy (anonymous):

just be more elaborate

OpenStudy (anonymous):

ok, have a great (night or day)

OpenStudy (anonymous):

thats not one of the options

OpenStudy (anonymous):

Lol, also u should close this question if u are done.